OpenStudy (anonymous):

Solve the equation for c and show your work. -2(c + 19) = -52

OpenStudy (anonymous):

-2(c + 19) = -52 divide the equation by -2 c + 19 = 26 c = 26 - 19 = 7

OpenStudy (anonymous):

-2(c+19)=-52 -2c-38=-52 +38 +38 -2c=-14 divide both sides by -2 c=7